# Enhance Eye Disease Detection using Learnable Probabilistic Discrete Latents in Machine Learning Architectures

### v2

@@ 4.1 Eye Disease Detection Experiment
We also performed experiments with noise added to the images, which revealed insightful results. Models equipped with GFlowOut showed enhanced performance compared to their standard counterparts, even under noisy conditions. Remarkably, the accuracy of these models with GFlowOut remained comparable to scenarios involving non-noisy data, underscoring the robustness of the model against different types of noise. This robustness is a significant finding, highlighting the model’s potential for practical applications where data imperfections are common.
Table 2: Experimental results of disease diagnosis . The above metrics mentioned are weighed averages. We note that the bottomup mask based on GFlowOut outperforms the other methods.
Precision Recall F1-Score Accuracy
ResNet18 none 0.66 0.58 0.61 52.72
random 0.70 0.64 0.66 55.50
bottomup 0.85 0.83 0.83 69.94
topdown 0.73 0.69 0.70 64.67
Vision Transformer none 0.64 0.68 0.65 69.04
random 0.70 0.66 0.67 75.52
bottomup 0.91 0.89 0.89 83.26
topdown 0.75 0.71 0.72 79.89
Table 3: Robustness to noise experiments, with Gaussian noise applied to the images. The above metrics mentioned are weighed averages. We note that the bottomup mask based on GFlowOut outperforms the other methods.
Precision Recall F1-Score Accuracy
ResNet18 none 0.64 0.58 0.60 49.72
random 0.67 0.66 0.65 52.66
bottomup 0.82 0.80 0.80 68.89
topdown 0.71 0.67 0.68 61.48
Vision Transformer none 0.62 0.67 0.64 68.62
random 0.69 0.66 0.67 71.65
bottomup 0.90 0.86 0.87 82.98
topdown 0.75 0.69 0.71 76.53
These results are in line with our expectations. The Vision Transformer, being both a larger and transformer-based model as compared to the ResNet-50 model, is expected to learn more features from the datasets and perform better at the task out of the models in consideration. Similarly, for a fixed backbone model, we expect the observed pattern in the various masks. The none mask performs the worst, since it is behaving as though there is no dropout. The random mask performs like a regular dropout layer, which is slightly better than having no dropout in these large models. topdown and bottomup perform better and the best respectively, since they take into consideration the previous layer’s context, and in the case of bottomup mask, the input data as well, to compute the probability distribution that is to be used for dropout.
@@ 4.2 Out of Distribution Evaluation and Entropy Calculations
Table 4: Robustness to noise experiments, with Salt noise applied to the images. The above metrics mentioned are weighed averages. We note that the bottomup mask based on GFlowOut outperforms the other methods.
Precision Recall F1-Score Accuracy
ResNet18 none 0.64 0.51 0.56 50.79
random 0.68 0.66 0.66 48.20
bottomup 0.80 0.81 0.80 67.50
topdown 0.73 0.64 0.68 62.15
Vision Transformer none 0.60 0.65 0.62 68.08
random 0.66 0.66 0.66 75.99
bottomup 0.84 0.88 0.85 79.44
topdown 0.70 0.68 0.68 77.13
Table 5: Robustness to noise experiments, with Speckle noise applied to the images. The above metrics mentioned are weighed averages. We note that the bottomup mask based on GFlowOut outperforms the other methods.
Precision Recall F1-Score Accuracy
ResNet18 none 0.58 0.59 0.58 51.22
random 0.69 0.60 0.64 52.22
bottomup 0.85 0.79 0.81 69.42
topdown 0.70 0.66 0.67 63.85
Vision Transformer none 0.62 0.65 0.63 67.15
random 0.69 0.68 0.68 75.36
bottomup 0.90 0.85 0.87 81.04
topdown 0.77 0.77 0.77 79.13
@@
To further explore the explainability of our model, we visualized the attention maps of the Vision Transformer model. Using the PyTorch GradCAM implementation Gildenblat & contributors (2021), we generated attention maps and overlaid these maps on the original sample images. This visualization highlights the regions of the image deemed important by the model, thereby enhancing our understanding of the model’s decision-making process.
Table 6: Entropy values on training and evaluation dataset
Dataset Min Entropy Max Entropy Avg Entropy
ODIR 0.408967 0.663598 0.506282
JSIEC 0.470288 0.69244 0.564921

### v3

@@ 4.1 Eye Disease Detection Experiment
We also performed experiments with noise added to the images, which revealed insightful results. Models equipped with GFlowOut showed enhanced performance compared to their standard counterparts, even under noisy conditions. We only report the results of the Vision Transformer model with bottomup mask, since this model performs the best on the standard datasets. Remarkably, the accuracy of these models with GFlowOut remained comparable to scenarios involving non-noisy data, underscoring the robustness of the model against different types of noise. This robustness is a significant finding, highlighting the model’s potential for practical applications where data imperfections are common.
Table 2: Experimental results of disease diagnosis on the ODIR dataset. The below metrics mentioned are weighed averages. We note that the bottomup mask based on GFlowOut outperforms the other methods. We also note that the GFlowOut methods perform better than the baseline, and the random mask performs nearly the same. This is expected because this mask acts as a normal dropout.
Precision Recall F1-Score Accuracy AUROC
ResNet18 benchmark 0.73 0.71 0.72 61.08 0.842
none 0.73 0.71 0.72 61.08 0.798
random 0.75 0.76 0.75 65.66 0.832
bottomup 0.80 0.78 0.80 68.82 0.864
topdown 0.77 0.69 0.73 66.67 0.844
Vision Transformer benchmark 0.72 0.58 0.64 72.04 0.883
none 0.73 0.74 0.74 67.04 0.811
random 0.79 0.79 0.79 72.52 0.846
bottomup 0.83 0.81 0.82 81.16 0.923
topdown 0.72 0.71 0.71 78.19 0.889
Table 3: Experimental results of disease diagnosis on the RFMiD dataset. The below metrics mentioned are weighed averages. We note that the bottomup mask based on GFlowOut outperforms the other methods. We also note that the GFlowOut methods perform better than the baseline.
Precision Recall F1-Score Accuracy AUROC
ResNet benchmark 0.94 0.91 0.93 89.1 0.947
none 0.90 0.88 0.89 87.8 0.932
random 0.93 0.91 0.92 89.4 0.951
bottomup 0.95 0.95 0.95 90.3 0.961
topdown 0.93 0.95 0.94 89.9 0.955
Vision Transformer benchmark 0.95 0.93 0.94 91.1 0.959
none 0.91 0.88 0.89 89.8 0.939
random 0.95 0.94 0.94 92.0 0.962
bottomup 0.96 0.95 0.96 93.1 0.969
topdown 0.96 0.94 0.95 92.6 0.964
Table 4: Robustness to noise experiments. We note that the model is quite robust to noise, which was added manually to both datasets. This is an important metric since in clinical scenarios, it is quite possible that the images acquired have some sort of noise in them.
Dataset Noise Precision Recall F1-Score Accuracy
ODIR Gaussian 0.81 0.81 0.81 80.94
Salt 0.76 0.72 0.74 75.24
Speckle 0.72 0.70 0.71 69.66
RFMiD Gaussian 0.92 0.91 0.91 91.3
Salt 0.90 0.87 0.88 90.8
Speckle 0.92 0.89 0.90 89.9
These results are in line with our expectations. The Vision Transformer, being both a larger and transformer-based model as compared to the ResNet-18 model, is expected to learn more features from the datasets and perform better at the task out of the models in consideration. Similarly, for a fixed backbone model, we expect the observed pattern in the various masks. The none mask performs the worst, since it is behaving as though there is no dropout. The random mask performs like a regular dropout layer, which is slightly better than having no dropout in these large models. topdown and bottomup perform better and the best respectively, since they take into consideration the previous layer’s context, and in the case of bottomup mask, the input data as well, to compute the probability distribution that is to be used for dropout.
@@ 4.2 Out of Distribution Evaluation and Entropy Calculations
To further explore the explainability of our model, we visualized the attention maps of the Vision Transformer model. Using the PyTorch GradCAM implementation (Gildenblat & contributors, 2021), we generated attention maps and overlaid these maps on the original sample images. This visualization highlights the regions of the image deemed important by the model, thereby enhancing our understanding of the model’s decision-making process.
Table 5: Entropy calculations. We use the random mask as our baseline, as that is similar to normal dropout, and compare it with the bottomup mask. The models trained on these datasets were evaluated against the JSIEC dataset. We note that our model has higher uncertainty on OOD dataset than the baseline model.
Dataset Mask Value
ODIR random $0.37\pm 0.01$
bottomup $0.50\pm 0.12$
RFMiD random $0.36\pm 0.05$
bottomup $0.41\pm 0.09$
@@
Table 6: ECE Calculations. We note that the GFlowOut model performs better on both ID and OOD datasets.
Dataset Mask ECE Score
ID random 0.10468232
bottomup 0.06402796
OOD random 0.23651563
bottomup 0.14311012
In Table 6, we present the Expected Calibration Error (ECE) scores for both In-Distribution (ID) and Out-of-Distribution (OOD) datasets, evaluated using the baseline and GFlowOut models. The ECE scores are computed with $M=10$ bins. A lower ECE score indicates better model calibration, reflecting a closer alignment between predicted probabilities and actual outcomes.
Our results demonstrate that the GFlowOut model consistently achieves lower ECE scores compared to the baseline model. This finding indicates that GFlowOut provides better-calibrated predictions on both ID and OOD datasets. The improvement in calibration is particularly notable for the OOD dataset, where the GFlowOut model outperforms the baseline model, emphasizing its robustness in handling data beyond the training distribution.
Furthermore, we observe a correlation between model uncertainty, as measured by entropy, and ECE scores. Specifically, scenarios with higher uncertainty tend to exhibit higher ECE scores. This relationship suggests that calibration performance is influenced by the model’s confidence, with less confident predictions typically being less calibrated.
@@ 5 Conclusion
Appendix A Datasets
A.1 ODIR
The ODIR-5K dataset, released by a Chinese team, is a multi-label classification dataset of fundus images. It was made available during the “Intelligent Eye” competition hosted by Peking University in 2019 and contains paired fundus images of the left and right eyes from 5000 patients, with labeled data for 3500 cases released for training.
Unlike other fundus datasets (like CHASE and DRIVE), ODIR-5K’s key distinction lies in providing paired data for both left and right eyes along with relevant descriptions. This setup allows for the data to be used both as 7000 individual cases or as 3500 paired cases for exploring consistencies or other aspects. ODIR-5K provides 8 categories of labels: Normal (N), Diabetes (D), Glaucoma (G), Cataract (C), Age-related Macular Degeneration (AMD), Hypertension (H), Myopia (M), and Other diseases/abnormalities (O). These categories encompass many common ophthalmic diseases and present an opportunity to explore the long-tail problem of datasets where some categories have much less data than others.
The significance of developing local datasets like ODIR-5K is profound for constructing medical diagnostic AI that is more suitable for the Chinese population, addressing biases that datasets can bring into AI research. By tailoring datasets to reflect the demographic and clinical profiles of the local population, AI systems can deliver more accurate and reliable diagnoses, ultimately enhancing patient care.
Table 7: Details about ODIR-5K dataset
Tag Training Set Off-Site Test Samples On-Site Test Samples Total
N 1135 161 324 1620
D 1131 162 323 1616
G 207 30 58 307
C 211 32 64 243
A 171 25 47 295
H 94 14 30 138
M 177 23 49 249
O 944 134 268 1346
A.2 RFMiD
The RFMiD 2.0 dataset is a multi-label classification dataset for fundus images. It is an updated version of the RFMiD (1.0 version), which was used in the RIADD Challenge at ISBI 2021. Released in 2023, the update primarily includes modifications to the label categories, refining the previously general ’Other’ category into specific rare diseases, thereby enhancing the quality of the labels. The dataset contains a total of 3,200 cases, and as of now, images and labels for all cases have been provided. With annotations for 45 different ocular diseases, RFMiD 2.0 holds the distinction of having the most disease categories among publicly available fundus datasets.
Fundus images, due to their similarity to natural images, have been used as a benchmark for the generalizability of methods in the medical field by researchers of natural images. This includes datasets such as CHASE, DRIVE, etc., which have been discussed in our previous articles. As the fundus dataset with the most annotated categories available publicly, RFMiD 2.0 is of value to both general multi-label classification researchers and those working on computer-aided diagnostic research related to vision. Given the large number of annotation categories in the dataset, some categories inevitably have very few samples, making this dataset also relevant for research into multi-label long-tail problems and issues related to limited samples in the medical field.
Figure 2: Class count of data points in the RFMiD Dataset
A.3 JSIEC
The JSIEC dataset consists of a total of 209,494 fundus images, covering 39 categories. This article introduces a subset of the JSIEC dataset, which includes 1,000 fundus images distributed across 39 categories. This dataset collects fundus images from 7 different data sources for the development and validation of deep learning algorithms. The primary datasets for training, validation, and testing come from the Picture Archiving and Communication System (PACS) of the Joint Shantou International Eye Center (JSIEC) in China, China’s Lifeline Express Diabetic Retinopathy Screening System (LEDRS), and the Eye Picture Archiving and Communication System (EyePACS) in the USA.
Millions of people worldwide are affected by fundus diseases such as Diabetic Retinopathy (DR), Age-related Macular Degeneration (AMD), Retinal Vein Occlusion (RVO), Retinal Artery Occlusion (RAO), Glaucoma, Retinal Detachment (RD), and fundus tumors. Among these, DR, AMD, and Glaucoma are the most common causes of vision impairment in most populations. Without accurate diagnosis and timely appropriate treatment, these fundus diseases can lead to irreversible blurring of vision, visual distortion, field defects, and even blindness. However, in rural and remote areas, especially in developing countries, there is a lack of ophthalmic services and ophthalmologists, making early detection and timely referral for treatment often inaccessible. Notably, fundus photography provides basic detection of these diseases and is available and affordable in most parts of the world. Non-professionals can handle fundus photographs and send them online to major ophthalmic institutions for follow-up. Artificial intelligence technology can be used to assist in diagnosis.
Table 8: Details about the JSIEC dataset
Disease No. of Images Disease No. of Images
Normal 38 CRVO 22
Tessellated 13 Yellow-white spots-flecks 29
Large Optic Cup 50 Cotton-wool spots 10
DR1 18 Vessel tortuosity 14
Possible glaucoma 13 Chorioretinal atrophy-coloboma 15
Optic atrophy 12 Preretinal hemorrhage 10
DR2 49 Fibrosis 10
DR3 39 Laser Spots 20
Sever hypertensive 15 Silicon oil in eye 19
Disc swelling and elevation 13 Blur fundus without PDR 111
Dragged Disc 10 Blur fundus with suspected PDR 45
Congenital disc abnormality 10 RAO 16
Retinitis pigmentosa 22 Rhegmatogenous RD 57
Bietti Crystalline Dystrophy 8 CSCR 14
Peripheral Retinal Degeneration and Break 14 VKH Disease 14
Myelinated Nerve Fiber 11 Maculopathy 74
Vitreous Particles 14 ERM 26
Fundus Neoplasm 8 MH 23
BRVO 44 Pathological Myopia 54
Massive Hard Exudates 13
Appendix B Images